OpenStudy (anonymous):

Need help Please I have to have this turned in NOW!!!! 1.Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

I just helped you with this question! If you were confused you should have said so instead of closing the question!

OpenStudy (anonymous):

OK we had 2 equations. x+y = 56 and y = 8 + x do you know how to solve a system of equations?

OpenStudy (anonymous):

srry i was having a conversation with a teacher and i have to have this done and completed

OpenStudy (anonymous):

sorta but not really

OpenStudy (anonymous):

I like comparison. In both equations, solve for y=. It's already done for you in the second. Then set the two equations equal to each other, eliminating the y and solve for x.

OpenStudy (anonymous):

so just 8 + x

OpenStudy (anonymous):

=56

OpenStudy (anonymous):

No because you have to solve for y in the first equation. You get x+y = 56 y = 56-x and y= 8 + x so 8+x = 56-x solve for x!

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

would i first minus eight to both sides?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x = 48 -x

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

Add x on both sides

OpenStudy (anonymous):

ok so this gives me 48

OpenStudy (anonymous):

alone

OpenStudy (anonymous):

i mean 2x = 48

OpenStudy (anonymous):

ok now divide by 2

OpenStudy (anonymous):

24

OpenStudy (anonymous):

can u help me with a few more

OpenStudy (anonymous):

I'll help you with 1 more but you should be starting to be able to recognize the pattern and figure out how to solve it by yourself

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2.Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770?

OpenStudy (anonymous):

Okay here the variables (x and y) are the number of adults and the number of students. So we can say x = number of adults y = number of students These are the 2 statements: 1. A total of 150 people went on the trip 2. they paid a total of $4770 Can you turn these into equations?

OpenStudy (anonymous):

x + y = 150 x45.00 + y28.50 = 4770

OpenStudy (anonymous):

????

OpenStudy (anonymous):

@BluFoot

OpenStudy (anonymous):

Sorry was afk for a sec. Yes that's correct. Now use substitution to solve. Solve for x or y in the first equation and plug it into the second equation

OpenStudy (anonymous):

im not sure how to do that

OpenStudy (anonymous):

cause first one we did comparison

OpenStudy (anonymous):

Let's solve for y in the first equation x+y = 150 y= 150-x Now plug it into the second 45x + 28.5y = 4770 45x + 28.5(150-x) = 4770 now expand and solve for x

OpenStudy (anonymous):

umm ok ive never seen that done before now though

OpenStudy (anonymous):

OK I have to go, good luck :)

OpenStudy (anonymous):

oh ok thanks